OpenStudy (anonymous):

Two points A and C have coordinates (1,3) and (7,7). a. Find the equation of the perpendicular bisector of AC b. B is a point on the y-axis which is equidistant from A and C and ABCD is a rhombus find point B and Point D c. find the area of the rhombus ABCD d. calculate the perpendicular distance of A from BC

OpenStudy (anonymous):

i have found point B (0,11) but need point D!

OpenStudy (anonymous):

DIstance of B to the perpendicular of AC = Distance of D from AC . And the lines AC and BD cuts each other perpendicularly . Area of rhombus is 0.5( AC * BD)

OpenStudy (anonymous):

how do you find point D then?

OpenStudy (anonymous):

Well you have to find the area of the rhombus and you need to to mulitply the two diagonal length together with half , and you can find the distance of B to the diagonal using pythagoras , the length of BD is twice of that .

OpenStudy (anonymous):

i cant find the area if i dont have all my points can I?

OpenStudy (anonymous):

a) here slope of AC is (7-3)/(7-1) =2/3 midpoint of AC is (1+7)/2,(7+3)/2=>(4,5) equation of line perpendicular bisector to ac is (y-5)= -3/2 (x-4) or 2y-10=-3x+12 or 3x+2y=22

OpenStudy (anonymous):

b) bcoz B is a pt equidistant from A and C hence B must lie on the perpendicular bisector of AC i.e B lies on 3x+2y=22 let pt B be (0,b)[as B lies on y-axis hence we have 3.0+2.b=22 or b=11 hence cordinate of B is (0,11)

OpenStudy (anonymous):

and what about the coordinate of D?

OpenStudy (anonymous):

let the cordinates of D be (p,q) hence midpoint of BD and AC must coincide as ABCD is a rhombus thus midpoint of BD must be (4,5) i.e (0+p)/2,(11+q)/2 is (4,5 or (p/2)=4 and (11+q)/2=5 solving we have p=8 and q=-1 hence D is (8,-1)

OpenStudy (anonymous):

AC=sqrt((7-1)^2 +(7-3)^2)=sqrt(52)=2*sqrt(13) BD=sqrt((8-0)^2+(11+1)^2)=sqrt(208)=4*sqrt(13) hence area of ABCD =1/2 *AC*BD =1/2 *2*sqrt(13)*4*sqrt(13) = 52sq units

OpenStudy (anonymous):

now eq of line joining B(0,11) and C(7,7) is y-11=-4/7 *x or 4x+7y=77 distance of A(1,3) from BC is | (4.1 +7.3 -77)/sqrt(65) | =52/sqrt(65)

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

welcome dear

OpenStudy (anonymous):

i hope u understood each step and would be able to solve similar problem u face

OpenStudy (anonymous):

i have come across this type of question alot and struggle everytime. this will definately help!